OpenStudy (anonymous):

Solve The System. -7x+5y=-34 -3x-2=2

OpenStudy (saifoo.khan):

is it -2y?

OpenStudy (saifoo.khan):

in the second row?

OpenStudy (mathlegend):

@brandonloves in the 2nd row is it supposed to be -2y?

OpenStudy (anonymous):

yes, sorry

OpenStudy (mathlegend):

In this case.. I would multiply the first equation by 2... because I want to cancel out the y. Then I would multiply the 2nd equation by 5.

OpenStudy (mathlegend):

@brandonloves what do you get?

OpenStudy (anonymous):

oh ok, lemme check

OpenStudy (anonymous):

I got 2 for x, and I also need to find y, so I would multiply them again by -3 and 7 to get rid of the x value?

OpenStudy (mathlegend):

You would plug in 2.. into either equation.

OpenStudy (mathlegend):

for x of course

OpenStudy (mathlegend):

-7(2)+5y=-34 OR -3(2)-2y=2 solve one

OpenStudy (anonymous):

-14+5y=-34 and then move the -14 over, change the sign, divide the 5y and get y=-4

OpenStudy (anonymous):

or -6-2y=2 which would also be -4

OpenStudy (anonymous):

you rule, thanks!